AgI is an insoluble compound. Find the volume of a 0.0875M CaI2 solution required to eliminate Ag+ ions from 15.00 mL of a 0.0153M AgNO3 solution according to the reaction 2AgNO3(aq) + CaI2(aq) = 2AgI(s) + Ca(NO3)2(aq).

(A) 5.25 mL (B) 3.75 mL (C) 2.62 mL (D) 1.88 mL (E) 1.31 mL
Answer #1

Balanced chemical equation is:

2 AgNO3 + CaI2 ---> 2 AgI + Ca(NO3)2

Here:

M(AgNO3)=0.0153 M

M(CaI2)=0.0875 M

V(AgNO3)=15.0 mL

According to balanced reaction:

1*number of mol of AgNO3 =2*number of mol of CaI2

1*M(AgNO3)*V(AgNO3) =2*M(CaI2)*V(CaI2)

1*0.0153 M *15.0 mL = 2*0.0875M *V(CaI2)

V(CaI2) = 1.3114 mL

Answer: E